OpenStudy (anonymous):

Can someone explain how i do this - Bus math Given the net income, the percent budgeted for three variable expenses, and the amount of an unexpected bill, determine how much money should be taken from the indicated budget category to cover this bill. Remember to use the formula: item percent ÷ sum of the percents × amount of unexpected expense = amount taken from each budget category Net income: $279 Percent item 1: 25% Percent item 2: 10% Percent item 3: 7% Bill due: $78.50 Amount taken from item 3: __________

OpenStudy (anonymous):

Oh hi again!

jimthompson5910 (jim_thompson5910):

hello lol, one sec

jimthompson5910 (jim_thompson5910):

sum of percents: 25+10+7 = 42 so it looks like for item 3, you would get item percent ÷ sum of the percents × amount of unexpected expense = (7/42) * 78.50 = 13.083333333 = 13.08 so about 13.08 dollars

OpenStudy (anonymous):

what is the amount of unexpected expense, there is not one right?

jimthompson5910 (jim_thompson5910):

that's what I was hung up on for a sec

jimthompson5910 (jim_thompson5910):

but I think that's the Bill due: $78.50 line

jimthompson5910 (jim_thompson5910):

it's not the net income I don't think

OpenStudy (anonymous):

Oh ok!

jimthompson5910 (jim_thompson5910):

because it says Given the net income, the percent budgeted for three variable expenses, and the amount of an unexpected bill so the net income is already accounted for

OpenStudy (anonymous):

Also, if you plug in the net income it isn't one of the answers listed $25.75 , $13.08 $68.50 ) But 13.08 is! Thank you !

jimthompson5910 (jim_thompson5910):

ok great, glad some of my answers are lining up lol